# On an integrable 2+1-dimensional extended Dym equation: Lax pair, $\bar{\partial}$-dressing scheme and modulation

### v1

@@ Abstract
In 1+1-dimensions, an extension of the canonical solitonic Dym equation has previously been derived both in a geometric torsion evolution context and in the analysis of peakon solitonic phenomena in hydrodynamics. Here, a novel 2+1-dimensional S-integrable extended Dym-type equation is introduced. As Lax pair is constructed and an associated $\bar{\partial}$-dressing scheme detailed. Integrable modulated versions of the 2+1-dimensional extended Dym equation are generated via application of a class of involutory transformations with genesis in classical Ermakov theory.
1 INFN, Sezione di Lecce, Italy.
Email: konopelchenkob@gmail.com
2 School of Mathematics and Statistics,
University of New South Wales, Sydney, Australia.
Email: c.rogers@unsw.edu.au
3 Depto. de Matemática, Facultad de Ciencias Exactas y Naturales,
Universidad de Buenos Aires & IMAS-CONICET.
Ciudad Universitaria. Pabellón I (1428), Buenos Aires, Argentina
Email: pamster@dm.uba.ar
@@ 1 Introduction
Nonlinear moving boundary problems of Stefan-type for the extended Dym equation (1) have been shown in [8] to be amenable to exact solution via Painlevé II symmetry reduction and iterated application of a Bäcklund transformation. The procedure adopted involved linkage of this class of moving boundary problems to an associated exactly solvable class for the canonical solitonic Dym equation. Moving boundary problems for the latter and its reciprocal associates had previously been analysed in [9]. Reciprocal transformations linking the canonical AKNS and WK1 inverse scattering schemes have been set down in [10]. Therein, in addition, novel invariance of the Dym equation under a reciprocal transformation was established. This result was extended to the integrable Dym hierarchy in [11]. The reciprocal invariance admitted by the Dym equation, in principle, may be applied to generate additional associated classes of solvable nonlinear boundary problems for the extended Dym equation.
@@ 2 On Variants of the 2+1-Dimensional Dym Equation: Associated Linear Representations
admits a corresponding Lax pair
@@
It now proves convenient to rewrite (12) as the system
$\begin{array}[]{c}V_{t}=V^{3}(V_{x}-V_{xxx})+\dfrac{3}{V}(V^{2}\rho)_{y},\\[5.69054pt] V_{y}=-V^{2}\rho_{x}\end{array}$ (15)
@@
of (15)2 allows the introduction of the potential $W$ such that
@@
It is remarked that $W$ is analogous to potentials previously introduced in 2+1-dimensional soliton theory, notably for the canonical Kadomtsev-Petviasvili. The potential equation in $W$ is seen to be invariant under the scaling $x\rightarrow x^{\prime}=\lambda x,\lambda\in\mathbb{R}$.
@@ 3 Linear Representation Analysis
for some positive integer $N$. On substitution of the Laurent series (30) into (29) it is readily shown that $\mu$ necessarily has an essential singularity at $\lambda=0$. This impediment can be overcome by introduction of the ansatz
$F=\frac{1}{\lambda}f(x,y,t)+\frac{1}{\lambda^{2}}y+\frac{4i}{\lambda^{3}}t.$ (31)
@@ 4 The $\bar{\partial}$-Dressing Procedure
$L\chi(\lambda)\rightarrow 0\quad\text{as}\quad\lambda\rightarrow\infty.$ (37)
@@
which with $\phi=\chi e^{F}$ coincides with that in (16). Here, these equations admit solutions regular at $\lambda=0$, that is, with
@@
It is noted that the latter two equations contain additional terms $\dfrac{1}{f_{x}}$ in (45) and $\dfrac{1}{2f^{2}_{x}}$ in (46) respectively to the corresponding system obtained for the standard 2+1-dimensional Dym equation of [12]. Here, the pair (45), (46) completely characterizes the function $f(x,y,t)$ associated with the $\bar{\partial}$-dressing procedure for the 2+1-dimensional extended Dym equation (12).
It is of interest to record that (45), (46) together imply that
@@
which with $f=W$ is nothing but the 2+1-dimensional extended Dym potential equation (19). Indeed, (47) may be regarded as a novel alternative avatar of the 2+1-dimensional extended Dym equation as derived in connection with its associated $\bar{\partial}$-dressing procedure.
@@ 5.1 A Class of 2+1-Dimensional Involutory Transformations
In modern soliton theory, reciprocal-type Bäcklund transformations associated with admitted conservation laws were originally introduced in [24]. Conjugation was made therein with the classical Bianchi theorem of pseudo-spherical surface theory whereby multi-soliton solutions may be generated iteratively in an algorithmic manner Reciprocal transformations have subsequently proved to have extensive applications in soliton theory [qv [25], [26] and literature cited therein]. In nonlinear continuum mechanics, reduction to analytically tractable canonical systems via reciprocal transformations has been detailed in [27, 28]. In relativistic gasdynamics novel invariance under multi-parameter reciprocal transformations of canonical systems has recently been derived in [29, 30, 31].
@@ 5.2.2 Temporal Modulation
$u_{t}+2\partial_{x}(1-\partial_{xx})\left(\frac{1}{u^{1/2}}\right)+6u^{2}[\ u^{-1}\partial^{-1}_{x}(u^{1/2})_{y}\ ]_{y}=0$ (55)
@@ 5.3 Ermakov Modulation
Here, the constantly $c_{i}$, $i=1,2,3$ are such that
@@ 5.4 Ermakov-Painlevé Modulation
Hybrid Ermakov-Painlevé systems were originally derived in [38] via symmetry reduction of a multi-dimensional nonlinear Schrödinger system which incorporates a de Broglie-Bohm potential. Application was made therein via Ermakov Painlevé II symmetry reduction to the analysis of certain transverse wave motions in generalised Mooney-Rivlin hyperelastic materials. Ermakov-Painlevé II structure and integrable reduction has subsequently had physical application in such areas as cold plasma physics [39], Korteweg capillarity theory [40] and the and the analysis of Nernst-Planck electrolytic system boundary value problems [41].
Ermakov-Painlevé II-IV integrable modulation of coupled 1+1-dimensional solitonic systems of sine-Gordon, Demoulin and Manakov-type has been subsequently detailed in [42]. Integrable Ermakov-Painlevé modulation of established 2+1-dimensional solitonic systems can likewise be derived.

### v2

# Open Communications in Nonlinear Mathematical Physics ]ocnmp[ Vol.6 (2026) pp 1References Article

@@ Abstract
In 1+1-dimensions, an extension of the canonical solitonic Dym equation has previously been derived both in a geometric torsion evolution context and in the analysis of peakon solitonic phenomena in hydrodynamics. Here, a novel 2+1-dimensional S-integrable extended Dym-type equation is introduced. A Lax pair is constructed and an associated $\bar{\partial}$-dressing scheme detailed. Integrable modulated versions of the 2+1-dimensional extended Dym equation are generated via application of a class of involutory transformations with genesis in classical Ermakov theory.
@@ 1 Introduction
Nonlinear moving boundary problems of Stefan-type for the extended Dym equation (1) have been shown in [8] to be amenable to exact solution via Painlevé II symmetry reduction and iterated application of a Bäcklund transformation. The procedure adopted involved linkage of this class of moving boundary problems to an associated exactly solvable class for the canonical solitonic Dym equation. Moving boundary problems for the latter and its reciprocal associates has previously been analysed in [9]. Reciprocal transformations linking the canonical AKNS and WK1 inverse scattering schemes have been set down in [10]. Therein, in addition, novel invariance of the Dym equation under a reciprocal transformation was established. This result was extended to the integrable Dym hierarchy in [11]. The reciprocal invariance admitted by the Dym equation, in principle, may be applied to generate additional associated classes of solvable nonlinear boundary problems for the extended Dym equation.
@@ 2 On Variants of the 2+1-Dimensional Dym Equation: Associated Linear Representations
, admits a corresponding Lax pair
@@
It is now convenient to rewrite (12) as the system
$V_{t}=V^{3}(V_{x}-V_{xxx})+\dfrac{3}{V}(V^{2}\rho)_{y}$ (15a)
$V_{y}=-V^{2}\rho_{x}$ (15b)
@@
of (15a) allows the introduction of the potential $W$ such that
@@
It is remarked that $W$ is analogous to potentials previously introduced in 2+1-dimensional soliton theory, notably for the canonical Kadomtsev-Petviashvili equation. The potential equation in $W$ is seen to be invariant under the scaling $x\rightarrow x^{\prime}=\lambda x,\lambda\in\mathbb{R}$.
@@ 3 Linear Representation Analysis
for some positive integer $N$. On substitution of the Laurent series (30) into (29) it is readily shown that $\mu$ necessarily has an essential singularity at $\lambda=0$. This impediment can be overcome by introducing the ansatz
$F=\frac{1}{\lambda}f(x,y,t)+\frac{1}{\lambda^{2}}y+\frac{4i}{\lambda^{3}}t,$ (31)
where $f$ satisfies the necessary smoothness requirements.
@@ 4 The $\bar{\partial}$-Dressing Procedure
$L_{1}\chi(\lambda)\rightarrow 0\quad\text{as}\quad\lambda\rightarrow\infty.$ (37)
@@
which, with $\phi=\chi e^{F}$, coincides with that in (16). Here, these equations admit solutions regular at $\lambda=0$, that is, with
@@
It is noted that the latter two equations contain additional terms $\dfrac{1}{f_{x}}$ in (45) and $\dfrac{1}{2f^{2}_{x}}$ in (46), respectively, to the corresponding system obtained for the standard 2+1-dimensional Dym equation of [12]. Here, the pair (45), (46) completely characterizes the function $f(x,y,t)$ associated with the $\bar{\partial}$-dressing procedure for the 2+1-dimensional extended Dym equation (12).
It is of interest to note that (45), (46) together imply that
@@
which, with $f=W$, is nothing but the 2+1-dimensional extended Dym potential equation (19). Indeed, (47) may be regarded as a novel alternative avatar of the 2+1-dimensional extended Dym equation as derived in connection with its associated $\bar{\partial}$-dressing procedure.
@@ 5.1 A Class of 2+1-Dimensional Involutory Transformations
In modern soliton theory, reciprocal-type Bäcklund transformations associated with admitted conservation laws were originally introduced in [24]. Conjugation was made therein with the classical Bianchi theorem of pseudo-spherical surface theory whereby multi-soliton solutions may be generated iteratively in an algorithmic manner. Reciprocal transformations have subsequently proved to have extensive applications in soliton theory [qv [25], [26] and literature cited therein]. In nonlinear continuum mechanics, reduction to analytically tractable canonical systems via reciprocal transformations has been detailed in [27, 28]. In relativistic gasdynamics novel invariance under multi-parameter reciprocal transformations of canonical systems has recently been derived in [29, 30, 31].
@@ 5.2.2 Temporal Modulation
$u_{t}+2\partial_{x}(1-\partial_{xx})\left(\frac{1}{u^{1/2}}\right)+6u^{2}[\ u^{-1}\partial^{-1}_{x}(u^{1/2})_{y}\ ]_{y}=0,$ (55)
@@ 5.3 Ermakov Modulation
Here, the constants $c_{i}$, $i=1,2,3$ are such that
@@ 5.4 Ermakov-Painlevé Modulation
Hybrid Ermakov-Painlevé systems were originally derived in [38] via symmetry reduction of a multi-dimensional nonlinear Schrödinger system which incorporates a de Broglie-Bohm potential. Application was made therein via Ermakov Painlevé II symmetry reduction to the analysis of certain transverse wave motions in generalised Mooney-Rivlin hyperelastic materials. Ermakov-Painlevé II structure and integrable reduction has subsequently had physical application in such areas as cold plasma physics [39], Korteweg capillarity theory [40] and the analysis of Nernst-Planck electrolytic system boundary value problems [41].
Ermakov-Painlevé II-IV integrable modulation of coupled 1+1-dimensional solitonic systems of sine-Gordon, Demoulin and Manakov-type was subsequently detailed in [42]. Integrable Ermakov-Painlevé modulation of established 2+1-dimensional solitonic systems can likewise be derived.
Acknowledgements
We thank the anonymous reviewer for the careful reading of the manuscript and his/her fruitful suggestions.